OpenStudy (anonymous):

Divide (5x3 – 6 – 5x) ÷ (10 + 5x) using synthetic division.

OpenStudy (anonymous):

A. x2 – 2x + 3, R 24 B. x2 – 3x, R 44 C. x2 – 2x + 3, R –36 D. x2 – 3x, R –56

OpenStudy (anonymous):

you know synthetic division???

OpenStudy (anonymous):

no i dont .

OpenStudy (anonymous):

ok lets start

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you have this (5x3 – 6 – 5x) ÷ (10 + 5x) if i re arrange the numerator (5x3 – 6 – 5x) to this 5x3+0x2-5x-6 do you agree with this???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now write the coefficients of x3 , x2 , x and the constant which are 5 0 -5 6

OpenStudy (anonymous):

now consider the denominator (10 + 5x) put it equal to zero we get (10 + 5x) = 0 thus x = -10/5 = -2 => x= -2 clear up to here???

OpenStudy (anonymous):

if you are clear up to this point we can start the process of synthetic division

OpenStudy (anonymous):

howd you get -10/5

OpenStudy (anonymous):

look you have (10 + 5x) if you put it equal to zero you will get (10 + 5x) = 0 clear???

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

subtract 10 from both sides we get 10-10+5x = 0-10 => 5x = -10 clear??

OpenStudy (anonymous):

oh okay. then divide -10 by 5 and get -10/5.

OpenStudy (anonymous):

yes exactly and -10/5 results in -2

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thus x = -2

OpenStudy (anonymous):

now the actual process starts here

OpenStudy (anonymous):

take the coefficients 5 0 -5 6 and the rest is on the drawing

OpenStudy (anonymous):